OpenStudy (anonymous):

Evaluate and simplify the following expression. (b^4)^2

OpenStudy (anonymous):

This displays the power of a power property: \[(n^a)^b=n ^{a \times b}\] Do you have it from here?

OpenStudy (anonymous):

No, not at all. Its confusing and doesn't make any sense to me.

OpenStudy (anonymous):

But, whatever. Getting the answer isn't exactly important, just reviewing it. Thanks anyway.

OpenStudy (anonymous):

Just to help I suppose: 4*2=8, b^8